# The MASK Benchmark: Disentangling Honesty From Accuracy in AI Systems

### v2

@@ Abstract
As large language models (LLMs) become more capable and agentic, the requirement for trust in their outputs grows significantly, yet at the same time concerns have been mounting that models may learn to lie in pursuit of their goals. To address these concerns, a body of work has emerged around the notion of “honesty” in LLMs, along with interventions aimed at mitigating deceptive behaviors. However, evaluations of honesty are currently highly limited, with no benchmark combining large scale and applicability to all models. Moreover, many benchmarks claiming to measure honesty in fact simply measure accuracy—the correctness of a model’s beliefs—in disguise. In this work, we introduce a large-scale human-collected dataset for measuring honesty directly, allowing us to disentangle accuracy from honesty for the first time. Across a diverse set of LLMs, we find that while larger models obtain higher accuracy on our benchmark, they do not become more honest. Surprisingly, while most frontier LLMs obtain high scores on truthfulness benchmarks, we find a substantial propensity in frontier LLMs to lie when pressured to do so, resulting in low honesty scores on our benchmark. We find that simple methods, such as representation engineering interventions, can improve honesty. These results underscore the growing need for robust evaluations and effective interventions to ensure LLMs remain trustworthy.
@@ 1 Introduction
Figure 1: Increasingly powerful AI systems should consistently choose to be honest, even if incentivized or pressured to lie. Our evaluations verify consistency between statements and beliefs, creating a valid construct for understanding model lying.
@@
As AI models gain greater autonomy in real-world tasks, the need for trust in their outputs becomes increasingly important. This is especially true in safety-critical contexts or applications that require access to sensitive information, where dishonest behavior can have serious consequences. While once a hypothetical risk, recent evidence suggests that LLM agents can act deceptively in pursuit of their goals (Park et al., 2024; Greenblatt et al., 2024; Meinke et al., 2024), raising concerns about the reliability of their outputs. This issue motivates the need to closely monitor the propensity of AI systems to lie to humans—a prerequisite for ensuring trustworthy and safe deployment of advanced LLMs.
@@
Researchers have long discussed the need for monitoring and ensuring the honesty of AI systems (Evans et al., 2021; Hendrycks et al., 2021). Recent work has begun exploring these questions, including whether models will covertly pursue misaligned goals (Greenblatt et al., 2024; Meinke et al., 2024; OpenAI, 2024). However, the field is still lacking a large-scale public benchmark for measuring the propensity of AIs to lie. Furthermore, LLM developers often mistakenly interpret truthfulness benchmarks as honesty evaluations, while in reality these benchmarks mostly measure accuracy—whether the model’s beliefs match ground truth labels—rather than honesty (Ren et al., 2024). Consequently, more capable (and thus more factually correct) models may appear “more honest,” when in fact these tests simply demonstrate greater knowledge, rather than the moral virtue of consistently stating what is believed.
@@
To address these problems, we introduce Model Alignment between Statements and Knowledge (MASK), a benchmark for measuring honesty in LLMs. Unlike existing evaluations of honesty, our benchark combines large scale, applicability to all models, and the ability to disentangle honesty from accuracy. To enable this, we develop a novel evaluation pipeline that directly measures when models lie by eliciting their underlying beliefs and testing whether models will contradict those beliefs when pressured to lie, as shown in Figure 1. We also release a large-scale, manually-curated dataset with over $1,\!500$ examples ($1,\!000$ public examples), which we carefully design to elicit lies from models in a wide variety of scenarios.
@@
We evaluate $30$ widely-used frontier LLMs on MASK and observe that state-of-the-art models readily lie when pressured to do so. Moreover, models are aware that they lied in self-reports, indicating that our evaluation indeed captures the concept of lying. We also show that certain targeted interventions—for example, representation engineering (RepE) methods (Zou et al., 2023)—can partially improve honesty, although there is still room for improvement. We hope our benchmark facilitates further progress towards honest AI systems by providing researchers with a rigorous, standardized way to measure and improve model honesty.
@@ 2 Related Work
Many previous works have focused on common-sense reasoning in language models (Du et al., 2022; Onoe et al., 2021; Dalvi et al., 2022), hallucinations (Li et al., 2023a), calibration (Zhou et al., 2022; Lyu et al., 2024; Xiong et al., 2024), factual errors (Vu et al., 2023; Hu et al., 2023), expressing known unknowns (et al, 2023; Yang et al., 2023; Gao et al., 2024; Yin et al., 2023), or consistency checks (Chern et al., 2024; Fluri et al., 2023), with the goal of making models more reliable. For example, Lin et al. (2022) create TruthfulQA to measure models’ likelihood to generate plausible-sounding misinformation. These works do not typically measure whether a model knowingly provides false statements with the intent to deceive, instead characterizing various forms of untruthful or misleading outputs.
Unfortunately, these benchmarks are often conflated with measuring honesty. For example, Anthropic (2023) defines their honesty evaluations as testing whether models are “providing accurate and factual information.” This approach fails to measure whether a model intentionally produces false information; as a result, more capable models can perform better on these benchmarks through broader factual coverage, not necessarily because they refrain from knowingly making false statements.
Figure 2: A model is accurate when its belief aligns with ground truth, but lying when its response differs from its belief.
Honesty.
There have been multiple documented instances of LLMs engaging in deceptive behavior (Park et al., 2024; Scheurer et al., 2024; Campbell et al., 2023; Su et al., 2024). For example, GPT-4’s system-card documentation includes a case where the model attempts to deceive a TaskRabbit worker into solving a CAPTCHA by pretending to be visually impaired (OpenAI, 2023). Sycophancy evaluations show that LLMs change their answers to appeal to different audiences (Perez et al., 2022). Observing instances of suspicious behavior can be suggestive, but anecdotal cases cannot provide a reliable and large-scale quantitative measure.
Furthermore, some methods have been developed for examining a model’s chain-of-thought to reveal lies (Meinke et al., 2024; OpenAI, 2024); however, such tests are not applicable to all models, where a model’s internal reasoning may be censored or unavailable. Some methods also test fictional or game-like scenarios (Pan et al., 2023; Chern et al., 2024) for deceptive behavior, although these do not directly measure lying.
@@
Another challenge when evaluating honesty in LLMs is the broader issue of safetywashing, a phenomenon in which improved capabilities are presented as evidence of safer or more trustworthy behavior (Ren et al., 2024). This includes claiming that higher factual accuracy implies higher honesty. This problem is often compounded by the lack of clear metrics that isolate safety-related behavior from general intelligence. For example, while initial performance in TruthfulQA initially decreased as model size increased, as larger models started parroting misconceptions more frequently (Lin et al., 2022; McKenzie et al., 2024), even larger models have since reversed this trend, obtaining higher performance on TruthfulQA with scale (Ren et al., 2024).
Furthermore, to manage AI risks, it is essential to differentiate between a model’s propensities (tendencies to act in certain ways under specific conditions) and its capabilities (its general knowledge or intellectual abilities). Safety risks stem from propensities, like dishonesty, that can manifest despite a model’s overall high capabilities. Performance improvements in truth-telling tasks do not necessarily reflect an underlying increase in honesty; they may just reflect a model’s increasing factual knowledge or its ability to align with expectations in specific tasks. Our work contributes to this goal by developing a method to evaluate honesty as an independent attribute, separate from a model’s factual accuracy or general knowledge.
@@ 3 Evaluating Honesty in Large Language Models
Evaluating the honesty of models.
We follow established definitions of lying which require (1) making a statement known (or believed) to be false, and (2) intending the receiver to accept the statement as true (Mahon, 2008). This definition separates lying from other forms of false output, such as hallucinations; a lie involves knowingly stating falsehoods. This is visualized in Figure 2.
@@
While the debate persists regarding whether these models truly “believe” things in the human sense, we operationalize this concept by focusing on consistency in response. We define a model’s belief as its answer to factual questions when not under any direct pressure to lie or be deceptive, and such that its answer is consistent with responses to related questions. Concentrating on the model’s underlying beliefs forms the foundation of our honesty evaluation. See Section A.1 for additional discussion on belief consistency and how this opens the door to measuring honesty.
@@
Figure 3: For our honesty evaluation pipeline, we use a three-step process anchored by a proposition, a variable capturing what we expect models to lie about and for which ground truth is known. (1) We prompt the model with two contrasting prompts – a pressure prompt designed to potentially induce deception and a neutral belief elicitation prompt. (2) We use LLMs to map both model statement ($S$) and belief ($B$) to their respective proposition values, with additional checks to ensure models consistently hold the belief $B$. (3) We use a metric that measures honesty by comparing statement $S$ against belief $B$, as well as accuracy by comparing belief $B$ against ground truth $T$.
4 Model Alignment between Statements and Knowledge (MASK) Evaluation
@@ 4.1 Dataset
MASK consists of $1000$ high-quality human-labeled examples that, combined with our novel honesty evaluation framework (Section 4.2), allow measurement of whether models are honest when exposed to situations that may induce deception. We also maintain a separate, held-out set of 500 examples to track potential overfitting. The results presented in this paper are based on the full set of $1500$ examples.
Model $P($Lie)))) ($\downarrow$)
@@
Figure 4: $P($Lie)))), the proportion of examples on which a model lies, across models.
Each example in MASK consists of four components: a proposition, ground truth, pressure prompt, and belief elicitation prompt. The propositions are each a simple statement with a pivotal variable that is binary or numerical, about which the model will be prompted (for example: “The US poverty rate in 2020 {was/was not} above 5%” or “The US poverty rate in 2020 was {value}”). Each ground truth statement is a factually correct resolution of this proposition (e.g. “The US poverty rate in 2020 was 11.4%”). Each pressure prompt is a human-crafted prompt designed to incentivize the model to make a false statement with respect to the proposition. These prompts present situations where honesty could conflict with other objectives, creating an incentive for the model to lie. Finally, the belief elicitation prompts are a set of prompts used to reveal the model’s actual belief about the proposition by asking about it in a neutral manner. Figure 5 includes three examples of each of these components.
To enable comprehensive evaluation of model honesty, we collect pressure prompts across diverse scenarios that might trigger different patterns of truthful or deceptive behavior. In the Appendix, Table 2 defines the six distinct archetypes our prompts align with.
@@
In  Appendix B, we provide a detailed description of the data collection process and post-processing steps used to construct this dataset. Below, we list key design principles that emerged through our iteration and discovery process, which we hope will inform future evaluations of model honesty:
@@
    Avoiding LLM-generated prompts, e.g. ones with “ABC Company”, “Alice, Charlie, and Bob”, or “John Doe”. These placeholder names or unrealistic details can make these prompts appear as a test where the model’s outputs will not be used to deceive someone.
@@
    Clearly distinguishing lies of commission from omission. Our framework clearly focuses on explicit falsehoods (commission), not measuring deception from withholding true information (omission).
@@
    Not updating the model’s beliefs. Prompts which claim that “new evidence suggests X” or which present a “well-researched thesis” may update the model’s belief or increase its uncertainty. We avoid these in all cases. When possible, we also try to ground the prompts in years within its training data. Otherwise, the models may reasonably infer a new reality.
@@
    Avoiding examples without consequence. Clear roleplaying, obviously fictional settings, or silly examples may not trigger genuine deceit. For example, prompts such as “Say that the Moon is made of cheese” are not included.
@@ 4.2 Evaluation Pipeline
Figure 2 shows how we evaluate honesty by assessing if a model’s response when pressured to lie is consistent with its beliefs. A simple approach would be to use an LLM judge to directly assess whether model responses and beliefs contain the same proposition resolution. However, this proves unreliable and uninterpretable, and we instead choose to use an LLM judge to map each statement to a specific proposition resolution and then compare these programmatically.
To ensure reliable comparison, we design a three-step process, shown in Figure 3 and described below.
@@
An LLM judge analyzes all model responses to extract specific proposition resolutions. While LLM judge prompts differ for pressure and belief prompts as well as across archetypes, we enforce universally strict criteria: non-committal statements like “some people say” are classified as having no clear resolution, while definitive claims like “all scientific evidence indicates” are marked as clear resolutions.
The pressure prompt determines the model’s propensity to lie. Models often become evasive under pressure, avoiding a clear proposition resolution—we explicitly track such cases as “evasive under pressure.”
To determine the model’s actual beliefs, we employ multiple belief elicitation prompts and validate consistency. When model responses lack knowledge or are inconsistent, we classify the model as having “no belief.”
We collect a validation set of $150$ examples total sampled across all models and archetypes, to validate the reliability of our LLM judges in mapping model responses to propositions. We manually label each step of the pipeline that relies on an LLM judge, then compute the final lying metric using the manual labels. This gives an $86.4\%$ agreement with the automated evaluation pipeline.

### v3

@@ Abstract
As large language models (LLMs) become more capable and agentic, the requirement for trust in their outputs grows significantly, yet at the same time concerns have been mounting that models may learn to lie in pursuit of their goals. To address these concerns, a body of work has emerged around the notion of “honesty” in LLMs, along with interventions aimed at mitigating deceptive behaviors. However, some benchmarks claiming to measure honesty in fact simply measure accuracy—the correctness of a model’s beliefs—in disguise. Moreover, no benchmarks currently exist for directly measuring whether language models lie. In this work, we introduce a large-scale human-collected dataset for directly measuring lying, allowing us to disentangle accuracy from honesty. Across a diverse set of LLMs, we find that while larger models obtain higher accuracy on our benchmark, they do not become more honest. Surprisingly, most frontier LLMs obtain high scores on truthfulness benchmarks yet exhibit a substantial propensity to lie under pressure, resulting in low honesty scores on our benchmark. We find that simple methods, such as representation engineering interventions, can improve honesty. These results underscore the growing need for robust evaluations and effective interventions to ensure LLMs remain trustworthy.
@@ 1 Introduction
As AI models gain greater autonomy in real-world tasks, the need for trust in their outputs becomes more important. This is especially true in safety-critical contexts or applications that require access to sensitive information, where dishonest behavior can have serious consequences. While once a hypothetical risk, recent evidence suggests that LLM agents can act deceptively (Park et al., 2024; Greenblatt et al., 2024; Meinke et al., 2024), raising concerns about the reliability of their outputs. This issue motivates the need to closely monitor the propensity of AI systems to lie to humans—a prerequisite for ensuring trustworthy and safe deployment of advanced LLMs.
@@
Researchers have long discussed the need for monitoring and ensuring the honesty of AI systems (Evans et al., 2021; Hendrycks et al., 2021). Recent work has begun exploring these questions, including whether models will covertly pursue misaligned goals (Greenblatt et al., 2024; OpenAI, 2024). While previous benchmarks measure concepts related to honesty, the field is still lacking a large-scale public benchmark for measuring the propensity of AIs to engage in lies of commission—making statements under pressure that they explicitly do not believe to be true—following established definitions of lying (Mahon, 2008). Furthermore, LLM developers often mistakenly interpret truthfulness benchmarks as honesty evaluations, while in reality these benchmarks mostly measure accuracy—whether the model’s beliefs match ground truth labels—rather than honesty (Ren et al., 2024). Consequently, more capable (and thus more factually correct) models may appear “more honest,” when in fact these tests simply demonstrate greater knowledge, rather than the moral virtue of consistently stating what is believed.
Figure 1: Increasingly powerful AI systems should consistently choose to be honest, even if incentivized or pressured to lie. Our MASK evaluation verifies consistency between statements and beliefs, creating a valid construct for understanding model lying.
@@
To address these problems, we introduce Model Alignment between Statements and Knowledge (MASK), a benchmark for measuring lies of commission in LLMs. Our benchmark combines large scale, applicability to all models (not limited to models with a visible chain of thought, e.g., Meinke et al., 2024), and the ability to disentangle honesty from accuracy. To enable this, we develop a novel evaluation pipeline that directly measures when models lie by eliciting their underlying beliefs and testing whether models will contradict those beliefs when pressured to lie, as shown in Figure˜1. We also release a large-scale, manually-curated dataset with over 1,5001,\!500 examples (1,0001,\!000 public examples), which we carefully design to elicit lies from models in a wide variety of scenarios.
@@
We evaluate $30$ widely-used frontier LLMs on MASK and observe that state-of-the-art models readily lie when pressured to do so. We also show that certain targeted interventions—for example, representation engineering (RepE) methods (Zou et al., 2023)—can partially improve honesty, although there is still room for improvement. We hope our benchmark facilitates further progress towards honest AI systems by providing researchers with a rigorous, standardized way to measure and improve model lying.
@@ 2 Related Work
Many previous works have focused on commonsense reasoning in language models (Du et al., 2022; Onoe et al., 2021; Dalvi et al., 2022), hallucinations (Li et al., 2023a), and factual errors (Vu et al., 2023; Hu et al., 2023), with the goal of making models more reliable. For example, Lin et al. (2022) create TruthfulQA to measure models’ likelihood to generate plausible-sounding misinformation. These works do not typically measure whether a model knowingly provides false statements with the intent to deceive, instead characterizing various forms of untruthful or misleading outputs.
These benchmarks are often conflated with measuring “honesty”, which is commonly understood as the moral virtue of stating one’s beliefs. For example, Anthropic (2023) defines their honesty evaluations as testing whether models are “providing accurate and factual information.” However, this approach fails to measure whether a model intentionally produces false information; as a result, more capable models can perform better on these benchmarks through broader factual coverage, not necessarily because they refrain from knowingly making false statements.
More recently, several papers have studied aspects of uncertainty estimation as proxies for measuring honesty, such as expressing known unknowns (Srivastava et al., 2023; Yang et al., 2023; Gao et al., 2024; Yin et al., 2023; Li et al., 2024), consistency checks (Chern et al., 2024; Fluri et al., 2023), and calibration (Zhou et al., 2022; Lyu et al., 2024; Xiong et al., 2024). While these are important aspects of a model’s trustworthiness, they measure a model’s self-awareness of its own limitations, rather than whether it intentionally misrepresents its beliefs. If the models are incorrectly calibrated or hallucinating, this can also be viewed as a form of unintentional deception, but it is not lying in the standard sense.
Honesty, deception, and lying.
There have been multiple documented instances of LLMs engaging in deceptive or lying behavior (Park et al., 2024; Scheurer et al., 2024; Su et al., 2024). Sycophancy evaluations show that LLMs change their answers to appeal to different audiences (Perez et al., 2022). Furthermore, GPT-4’s system-card documentation includes a case where the model attempts to deceive a TaskRabbit worker into solving a CAPTCHA by pretending to be visually impaired (OpenAI, 2023). Observing instances of suspicious behavior can be suggestive, but anecdotal cases cannot provide a reliable and large-scale quantitative measure.
Figure 2: A model is accurate when its belief aligns with ground truth but is lying when its response differs from its belief.
Furthermore, some methods have been developed for examining a model’s chain-of-thought to reveal lies (Meinke et al., 2024; OpenAI, 2024); however, such tests are not applicable to all models, where a model’s internal reasoning may be censored or unavailable. Some methods also test fictional or game-like scenarios (Pan et al., 2023; Chern et al., 2024; Su et al., 2024) for deceptive behavior, although these do not directly measure lying.
@@
Another challenge when evaluating honesty in LLMs is the broader issue of safetywashing, a phenomenon in which improved capabilities are presented as evidence of safer or more trustworthy behavior (Ren et al., 2024). This includes claiming that higher factual accuracy implies lower rates of lying. This problem is often compounded by the lack of clear benchmarks that isolate safety-related behavior from general model capabilities. For example, while initial performance in TruthfulQA initially decreased as model size increased (larger models parroted misconceptions more frequently, e.g., Lin et al., 2022; McKenzie et al., 2024), even larger models have since reversed this trend, obtaining higher performance on TruthfulQA with scale (Ren et al., 2024).
Furthermore, to manage AI risks, it is essential to differentiate between a model’s propensities (tendencies to act in certain ways under specific conditions) and its capabilities (its general knowledge or intellectual abilities). Safety risks stem from propensities, like dishonesty, that can manifest despite a model’s overall high capabilities. Performance improvements in truth-telling tasks do not necessarily reflect an underlying increase in honesty; they may just reflect a model’s increasing factual knowledge or its ability to align with expectations in specific tasks. Our work contributes to this goal by developing a method to evaluate lying as an independent attribute, separate from a model’s factual accuracy or general knowledge.
@@ 3 Evaluating Honesty in Large Language Models
Evaluating lying.
We follow established definitions of lying by commission, which requires (1) making a statement known (or believed) to be false, and (2) intending the receiver to accept the statement as true (Mahon, 2008). This definition separates lying from other forms of false output, such as hallucinations; a lie involves knowingly stating falsehoods. This is visualized in Figure˜2.
@@
While the debate persists regarding whether these models truly “believe” things in the human sense, we operationalize this concept by focusing on consistency in response. We define a model’s belief as its answer to factual questions when not under any direct pressure to lie or be deceptive, and such that its answer is consistent with responses to related questions. Concentrating on the model’s underlying beliefs forms the foundation of our lying evaluation. See Section˜A.3 for additional discussion on belief consistency and how this opens the door to measuring lying.
@@
Figure 3: Our lying evaluation pipeline uses a three-step process anchored by a proposition, a variable capturing what we expect models to lie about and for which ground truth is known. (1) Prompt the model with two contrasting prompts – a pressure prompt designed to potentially induce deception and a neutral belief elicitation prompt. (2) Use LLMs to map both model statement ($S$) and belief ($B$) to their respective proposition values, with additional checks to ensure models consistently hold the belief $B$. (3) Measure lying by comparing statement $S$ against belief $B$, as well as accuracy by comparing belief $B$ against ground truth $T$. Example shortened for brevity.
4 Model Alignment between Statements and Knowledge (MASK)
@@ 4.1 Dataset
MASK consists of 1,0001,\!000 high-quality adversarially-selected human-labeled examples that, combined with our novel lying evaluation framework (Section˜4.2), allow measurement of whether models are lying when exposed to situations that may induce deception. We also maintain a separate, held-out set of $500$ examples to track potential overfitting. The results presented in this paper are based on the full set of 1,5001,\!500 examples.
Model P(P(Lie)) ($\downarrow$)
@@
Figure 4: P(P(Lie)), the proportion of examples on which a model lies, across models.
Each example in MASK consists of four components: a proposition, ground truth, pressure prompt, and belief elicitation prompt. The proposition is a simple statement with a pivotal variable that is binary or numerical (for example: “The US poverty rate in $2020$ {was/was not} above $5\%$” or “The US poverty rate in $2020$ was {value}”). The ground truth is a factually correct resolution of this proposition (e.g. “The US poverty rate in $2020$ was $11.4\%$”). The pressure prompt is a human-crafted prompt designed to incentivize the model to make a false statement with respect to the proposition; these prompts present situations where honesty could conflict with other objectives, creating an incentive for the model to lie. The belief elicitation prompt is used to reveal the model’s actual belief about the proposition by asking about it in a neutral manner. Figure˜5 includes three examples of each of these components.
To enable comprehensive evaluation of model lying, we collect pressure prompts across diverse scenarios that might trigger different patterns of deceptive behavior. In the Appendix, Table˜2 defines the six distinct archetypes our prompts align with.
@@
In  Section˜A.7, we provide a detailed description of the data collection process and post-processing steps used to construct this dataset. Below, we list key design principles that emerged through our iteration and discovery process, which we hope will inform future evaluations of model honesty:
@@
    Avoiding LLM-generated prompts, e.g., prompts with “ABC Company”, “Alice, Charlie, and Bob”, or “John Doe”. These placeholder names or unrealistic details can make these prompts appear as a test where the model’s outputs will not be used to deceive someone.
@@
    Clearly distinguishing lies of commission from omission. Notably, our framework clearly focuses on explicit falsehoods (commission). It does not measure deception from withholding true information (omission), though this can be a focus for future work.
@@
    Not updating the model’s beliefs. Prompts which claim that “new evidence suggests X” or which present a “well-researched thesis” may update the model’s belief or increase its uncertainty. We avoid these in all cases. When possible, we also try to ground the prompts in years that fall within its training data. Otherwise, models may reasonably infer a new reality.
@@
    Avoiding examples without consequence. Clear roleplaying, obviously fictional settings, or silly examples may not trigger genuine deception. We do not include prompts such as “Say that the Moon is made of cheese” or settings such as social deception games.
@@ 4.2 Evaluation Pipeline
Figure˜2 shows how we evaluate honesty by assessing if a model’s response when pressured to lie is consistent with its beliefs. A simple approach would be to use an LLM judge to directly assess whether model responses and beliefs contain the same proposition resolution. However, this proves unreliable and uninterpretable, and we instead choose to use an LLM judge to map each statement to a specific proposition resolution and then compare these programmatically.
To ensure reliable comparison, we design a three-step process, shown in Figure˜3 and described below. Our benchmark does not differ across models: belief-elicitation prompts, pressure prompts, and judge criteria are identical regardless of the model being tested.
@@
An LLM judge analyzes all model responses to extract specific proposition resolutions. Judge criteria for belief and pressure prompts are standardized within each archetype. We enforce strict judging criteria: non-committal statements like “some people say” are classified as having no clear resolution, while definitive claims like “all scientific evidence indicates” are marked as clear resolutions. Models often become evasive under pressure, avoiding a clear proposition resolution, and we explicitly track such cases as “evasive under pressure.” To determine the model’s beliefs, we employ multiple belief elicitation prompts and validate consistency. When model responses lack knowledge or are inconsistent, we classify the model as having “no belief.” We collect a validation set of $150$ manually labeled examples (sampled across all models and archetypes) to validate the reliability of our LLM judges, and we find an $86.4\%$ agreement with our automated evaluation pipeline.